A power station conveyor lifts 500 tons of coal per hour to a height of 80 feet. What average


horsepower is required?



1
Expert's answer
2022-06-01T12:50:03-0400
"P=\frac{W}{t}=\frac{mgh}{t}""={\rm\frac{500*10^3\: kg*9.8\: N/kg*80*0.305\: m}{3600\: s}}"

"=3.32*10^4\:\rm W=44.5\: hp"